record the main scale reading and vernier coincidence.
calculate the diameter using formula Diameter= main scale reading+(vernier coincidencex LC) - ZE.
LC= 0.05
ZE= 0
ungoro ....?.F.1.8.7...? பாப்ப்ப்ப்ப்ப்ப்ப்பும் | llll
Answer #1

From the figure, the vernier coincidence is at 7.5 and thus the vernier scale division= 2*7.5=15

And the main scale reading is 17mm.

And LC= 0.05

Thus, the diameter = main scale reading+(vernier coincidence×LC) - ZE = 17 + (15×0.05) - 0 mm = 17.75 mm